OpenStudy (loser66):

For each question on a multiple choice test, there are 5 possible answers, of which exactly one is correct. If a student selects answers at random, give the probability that the first question answer correctly is question 4. Please, help. I don't know how to start but \(P(x_i)=\dfrac{1}{5}\)

OpenStudy (loser66):

If the first correct answer is question 4, then the first 3 trial is failed so that P(1') =4/5 P(2') =4/5 P(3')=4/5 next is P(4) =1/5 then??

OpenStudy (kamille):

I think you need to multiply P(1'), P(2') P(3') P(4')

OpenStudy (loser66):

I know, but what I need is how to put it in logic. I can just give out the answer :)

OpenStudy (loser66):

So, the probability to get the 4th one is the first correct answer is P(first failed + second failed+third failed + fourth win) = \[P(1'\cap 2'\cap 3'\cap 4)=P(1')*P(2')*P(3')*P(4)=(\dfrac{4}{5})^3*\dfrac{1}{4}\]

OpenStudy (loser66):

Am I right?

OpenStudy (kamille):

Hm. so he could do: 1) Answer first one, the probability is 1/5 2) Don't answer first one, answer second one, probability is 4/5*1/5 3) don't answer first one,don't answer second one, answer third one, probability is 4/5*4/5*1/5 4) Don't answer first one, dont answer the second one, don't answer the third one,answer the fourth one, probability is: 4/5*4/5*4/5*1/5 1) 1/5 2)4/20 3)16/125 4)64/625 Now all you need to do is sum up all the probabilities,I think

OpenStudy (loser66):

Thank you very much. I got you, just correct a little bit by replace "Don't answer" by " wrong answer" :)

OpenStudy (kamille):

sorry for my English, I am not a native speaker;) And you are welcome,I hope I am correct.. Maybe @waterineyes could check?

OpenStudy (anonymous):

He he he he, you just scared me..!!

OpenStudy (anonymous):

I got notification like : "Sorry for my english, I am not a native speaker.." and you tagged me.. I thought you found out, I don't know how, that my english is also very poor.. :)

OpenStudy (kamille):

haha,sorry about that! Your English seems to be alright;)

OpenStudy (anonymous):

I have got rusty on these topics.. :(